A small toy gun fires light foam projectiles by compressing a spring. As the spring relaxes it pushes the projectile.




The toy gun has a spring with a spring constant of 484 N m−1, the spring is compressed by 5.00 cm, and it fires a 10.0 g projectile.




At what speed does the projectile leave the toy gun? (in m s−1 to 3 s.f)

1
Expert's answer
2022-04-11T12:04:31-0400

"\frac{kx^2}2=\frac{mv^2}2,"

"v=x\sqrt\frac km =11~\frac ms."